OpenStudy (vera_ewing):

Which ratio would be used to determine how many moles of water are produced by reacting 15 mol of oxygen in excess hydrogen? 2H2 + O2 ----> 2H2O

OpenStudy (vera_ewing):

A. 1 mol O2/1 mol H2O B. 1 mol H2O/1 mol O2 C. 2 mol H2O/1 mol O2 D. 1 mol O2/2 mol H2O

OpenStudy (vera_ewing):

@ghuczek please help me

OpenStudy (vera_ewing):

@Somy help please

OpenStudy (anonymous):

In the balanced equation hydrogen and water are in a 1 : 1 ratio.

OpenStudy (vera_ewing):

so it's A?

OpenStudy (anonymous):

\[\color{green}{\checkmark}\]

OpenStudy (vera_ewing):

How many moles of lithium are needed to completely react with 4.6 mol Br2? 2Li + Br2 2LiBr A. 9.2 mol B. 8.4 mol C. 4.6 mol D. 2.3 mol

OpenStudy (vera_ewing):

What's the answer? @ghuczek

OpenStudy (anonymous):

Use ratios and proportions. Notice in the balanced equation that 2 moles of Li are needed for every mole of bromine. So,\[\color{green}{\dfrac{2~\text{mol of Li}}{1~ \text{mol of Br}_2} = \dfrac{x~\text{mol of Li}}{4.6~\text{mol of Br}_2}}\] Cross multiply:\[\color{green}{(2~\text{mol of Li}) \times (4.6~\text{mol of Br}_2) = (x~\text{mol of Li}) \times (1~\text{mol of Br}_2)}\] and solve for x.